A golfer recorded the following scores in 25 recent games:
88, 89, 102, 93, 99, 95, 94, 96, 84, 91, 95, 86, 95, 92, 102, 84, 104, 108, 92, 93, 92, 97, 96, 95, 82
a) Determine the mean and standard deviation
b) Assuming a normally distributed set of scores, what is the probability that the golfer will qualify for the club championship, which requires him to get a score under 83 on his next game?
1
Expert's answer
2020-05-18T20:09:42-0400

"82,84,84,86,88,89,91,92,92,92,93,93,94,"

"95,95,95,95,96,96,97,99,102,102,104,108"

a)


"mean=\bar{X}={1\over 25}(82+84+84+86+88+89+91+""92+92+92+93+93+94+95+95+95+95""+96+96+97+99+102+102+104+108)=93.76"

"\displaystyle\sum_{i=1}^{25}(X_i-\bar{X})^2=(82-93.76)^2+2(84-93.76)^2+""+(86-93.76)^2+(88-93.76)^2+(89-93.76)^2"

"+(91-93.76)^2+3(92-93.76)^2+2(93-93.76)^2+"

"+(94-93.76)^2+4(95-93.76)^2+2(96-93.76)^2+"

"+(97-93.76)^2+(99-93.76)^2+2(102-93.76)^2+"

"+(104-93.76)^2+(108-93.76)^2=960.56"

"s^2={\displaystyle\sum_{i=1}^{25}(X_i-\bar{X})^2\over n-1}={960.56\over 25-1}={120.07\over 3}\approx40.023333"

"s=\sqrt{s^2}=\sqrt{{120.07\over 3}}\approx6.3264"

b)

"X\sim N(\mu, \sigma^2)." Then


"Z={X-\mu\over \sigma}\sim N(0, 1)"

"\mu=93.76, \sigma=6.3264"


"P(X>83)=1-P(X\leq 83)=1-P(Z\leq {83-93.76\over 6.3264})\approx"

"\approx1-P(Z\leq -1.7008)\approx0.9554"